OpenStudy (anonymous):

5.Which of the following functions match these outcomes: f(1) = 6 f(2) = 8 f(3) = 10 f(4) = 12 f(x) = 5x + 1 f(x) = 4x + 2 f(x) = x2 + 4 f(x) = 2x + 4 f(x) = 12 - 2x

OpenStudy (anonymous):

the value inside the parenthesis are the values that you have to plug in the given equations :)

OpenStudy (anonymous):

for example f(1) = 5x +1 = 5(1)+1 = 6

OpenStudy (anonymous):

When I did it it fit in to all of them but it said i was wrong.

OpenStudy (anonymous):

maybe you should check it carefully and plug in all the values of your of x one by one so that you know the possible outcomes of the given equaions :)

OpenStudy (anonymous):

for example in equation 1 you can only plug in x= 1 because if you put other values of x then the outcome would not match the other values.

OpenStudy (anonymous):

Thanks I got it!

OpenStudy (anonymous):

no problem :D